# Improving How Agents Cooperate: Attention Schemas in Artificial Neural Networks

### v1

@@ Abstract
Growing evidence suggests that the brain uses an “attention schema” to monitor, predict, and help control attention. It has also been suggested that an attention schema improves social intelligence by allowing one person to better predict another. Given their potential advantages, attention schemas have been increasingly tested in machine learning. Here we test small deep learning networks to determine how the addition of an attention schema may affect performance on a range of tasks. First, we found that an agent with an attention schema is better at judging or categorizing the attention states of other agents. Second, we found that an agent with an attention schema develops a pattern of attention that is easier for other agents to judge and categorize. Third, we found that in a joint task where two agents paint a scene together and must predict each other’s behavior for best performance, adding an attention schema improves that performance. Finally, we find that the performance improvements caused by an attention schema are not a non-specific result of an increase in network complexity. Not all performance, on all tasks, is improved. Instead, improvement is specific to “social” tasks involving judging, categorizing, or predicting the attention of other agents. These results suggest that an attention schema may be useful in machine learning for improving cooperativity and social behavior.
@@ Introduction
Growing evidence suggests that the brain uses an “attention schema” to monitor, predict, and help control attention [1-5]. The attention schema may be used to model more than one’s own attention. It may also be directed to model the attention of others, affording better theory of mind, better predictions of the behavior of others, and therefore more effective social interaction [1,6-12]. In the present study, we ask whether adding an attention schema to an artificial agent improves its ability to make judgements about the attention of other agents, improves its ability to have its own attention judged by others, and improves its ability to coordinate with another agent in a simple joint task.
@@
Recently, versions of attention schemas have been incorporated into artificial neural networks to test whether the same principles apparent in the human brain might provide benefits in machine learning, and to determine whether lessons from the more controlled, artificial case might provide conceptual insights into human cognition [13-17]. Two studies in particular suggest that self-models may improve performance in cooperative tasks. In one study, Liu et al. [15] trained interacting groups of networks on visual tasks that required the agents to coordinate with each other. A transformer-style visual attention was used. In some conditions, networks were provided with an attention schema, in the form of a learned ability to predict the state of their own attention mechanism. With an attention schema, agents performed the social coordination tasks better. One interpretation is that when an agent is trained to predict its own attention states, it also becomes better at predicting the attention states of other agents in the same environment. In that hypothesis, learning to self-model leads to an improvement in the ability to model others and thus to an improvement in group task performance.
@@
In a second recent study on the effects of self-modeling, Premakurma et al. [16] trained artificial networks on a variety of discrimination tasks, including both visual and linguistic tasks. In some conditions (the self-model conditions), the agents were also trained on an auxiliary task in which they were rewarded for predicting the activation states of their own inner layers. The networks that were trained on the auxiliary task showed a consistent change. They learned to make themselves simpler, more regularized, and more parameter-efficient as they performed the primary task. These gains were found across all types of networks studied. It was hypothesized that when a network is given the auxiliary task of predicting its own internal activity states, one way for it to ensure better performance is to learn to make its internal states simpler and more regular. The simpler the internal states of the network, the easier they are to predict. Therefore, as the network learns to predict itself, it learns to simplify itself. According to that hypothesis, learning a self-model restructures the self, regularizing the network, and making internal states an easier target of predictive modeling.
@@
Based on these prior findings, we proposed the following two hypotheses about the relationship between self-models and social behavior. First, when an agent learns to self-model, such as when it learns an attention schema, it can transfer some of that learning toward predictively modeling other agents. Predictive modeling of other agents then becomes a useful part of interacting with them. Second, when an agent learns to self-model, it restructures itself to be more regularized and more predictable, and therefore the agent becomes a more effective target for being predictively modeled by other agents. Both changes – becoming better at modeling others, and becoming a more transparent target for being modeled by others – should render the agent a more effective member of a cooperative group.
@@
The purpose of the present study was to test these hypotheses about how self-modeling, and in particular attention schemas, might influence multi-agent interaction. We conducted three related experiments. Experiment 1 asked: Do attention schemas improve the ability of networks to model each other’s attention states? Experiment 2 tested an alternative explanation: Are the performance benefits of an attention schema merely a non-specific effect of added network complexity? Finally, Experiment 3 addressed the question of multi-agent cooperation: Do agents with attention schemas perform better on a simple joint task?
@@
The purpose of Experiment 1 was to train artificial agents with and without attention schemas and to test two hypotheses. Hypothesis 1 was that an agent with an attention schema will be better at learning to make categorical judgements about the attention of other agents. Hypothesis 2 was that when an agent has an attention schema, other agents will perform better at learning to make categorical judgements about its attention.
@@
Agents were first trained on a visual categorization task (e.g. categorize a picture as showing a golf ball or a garbage truck). We then froze the weights in all layers except a final block, and trained the agents on a new task, the “attention judgement” task. The task allowed us to test the ability of one agent to make categorical determinations about the attention states of another agent. The attention judgment task was modified from Ziman et al. [18], who demonstrated the ability of people to assess the attention states of others. In the human version of the task, participants looked at a spotlight of attention moving around a picture and judged whether the spotlight represented real attention (based on eye movements recorded previously from a real person), or whether it represented a randomly scrambled version of attention (the recorded attention locations presented in a random order). Since people were able to perform above chance, the human brain must contain information about the typical or predicted patterns of attention and must be able to recognize when those patterns have been violated. Thus, humans demonstrate an implicit predictive model of attention.
@@
To modify the task for the present study, we assigned agents to one of two roles: either the receiver or the sender. If the agent was assigned the sender role, it performed its originally trained visual categorization task. The attention activation values, for each trial of the task, were saved in the form of a multidimensional tensor. In one version, the tensor was left in its correct, original form. In another version, the attention values were randomly scrambled along one dimension of the tensor.
@@
If the agent was assigned the receiver role, it was first trained on the image categorization task; then its weights were frozen except in the final block of layers; and then it was trained on the attention judgement task. On each trial of the attention judgement task, instead of receiving a picture as an input, the agent was provided with an attention tensor; and instead of being rewarded for categorizing the picture, it was rewarded for categorizing the tensor as either a valid example of attention (recorded from the performance of a sender agent) or as an invalid example (constructed by randomly modifying a valid tensor). The task therefore tested the ability of the receiver agent to learn to recognize violations of predictable patterns in the attention behavior of the sender.
@@
The experiment used a 2$\times$2 design: the sender agent could be trained with or without an attention schema, and the receiver agent could be trained with or without an attention schema. We proposed the following hypotheses.
@@
Hypothesis 2: If the sender agent was trained with an attention schema, then the receiver agent will be able to learn the attention judgement task better. Such a result would suggest that when an agent contains an attention schema, it is, itself, an easier target for other agents to build predictive models of it.
@@
We trained networks on a visual categorization task. Given an image as input (for example, a picture of a golf ball on a lawn), the network was trained to arrive at a dichotomous categorization (for example, golf ball versus garbage truck). All networks and task environments were created using PyTorch [19] and all experiment code and data are publicly available at github.com/kathrynfarrell/attention_schemas_in_anns. We used an image-adapted implementation of the Attention Schema Neural Network architecture proposed by Liu et al. [15], which utilized a transformer-based attention mechanism. Figure 1 shows the architecture, with 1A showing the control architecture (control network: attention but no attention schema) and 1B showing the architecture that includes an added attention schema (attention schema network).
@@
Figure 1: Figure 1: The Attention-Schematic Neural Network (ASNN) architecture. A. The control model without an attention schema. B. The model with an attention schema, modified from Liu et al. [15]. The model is based around a standard vision transformer (ViT) block, which generates attention scores and applies them to the network’s representation of the image (Values), to produce a representation of the image with the attention applied (“applied attention”). The applied attention is then passed into a Recurrent Neural Network (RNN) module with learnable parameters. The output of this module (“$R_{out}$”) is a precursor to both the network’s final attention prediction and its dynamic selection of attention scores, both of which are produced by several linear layers with learnable parameters (the Predictive Layers and the Decision Control Layers, respectively). For the attention-revision computations, attention suppression or activation is achieved through processes beginning in the Decision Control Layers. Linear “suppressor” and “activator” layers in the Decision Control produce logit pairs, which are binarized using a Gumbel-Softmax distribution, then combined via element-wise multiplication. The binary mask resulting from that multiplication is then applied to the original attention scores (by element-wise multiplication) yielding the final attention—a sparser version of the original attention scores where some of the scores have been selectively “shut off” or reduced to zero. The final attention is then applied to the original image representation (Values), and the result is passed to the final MLP head, which classifies the image as belonging to a particular category.
@@
For details on how the attention schema mechanism works, see Liu et al. [15]. Briefly, the objectives of the schema network are twofold: learning to model and refine its own attention and learning to complete a task (e.g. classifying images). The schema network uses its applied attention in service of both goals: modeling its own attention (by predicting its final attention), and further editing the attention to its final refined state. To accomplish this, the schema network uses a transformer block with six attention heads, and passes the applied attention into a recurrent neural network module with learnable parameters. The output of this module ($R_{out}$) is further processed and used to model attention; specifically, it is passed into the Predictive Layers, which output the predicted attention. $R_{out}$ is also further processed and used to refine attention; it is passed into the Decision Control Layers which produce the final refined attention (with the Decision Control Layers employing activator units, suppressor units, and categorical reparameterization with Gumbel-Softmax). In this way, manipulations of the initial applied attention are used to generate both the prediction of the final attention as well as the final attention itself. Schema networks are trained on their combined ability to accurately complete the main task (e.g. classifying an image) and predict their own attention.
@@
To produce an attention tensor, we stacked the scores from the three attention heads into a 257 token $\times$ 257 token $\times$ 3 attention head tensor (257 tokens in the first two dimensions resulting from the 16$\times$16 image patches from the original input image plus one class token). From these tensors, we created the veridical and scrambled attention data that would be sent to the receiving networks for discrimination. The veridical attention tensors were created by reducing the size of the tensors from 257 $\times$ 257 $\times$ 3 to 256 $\times$ 256 $\times$ 3, matching the size of the original input image. For the scrambled attention, we randomly permuted each tensor along its last dimension, which randomly shuffled the attention scores within the data from each attention head. Then, we reduced the resulting tensor to 256 $\times$ 256 $\times$ 3, again matching the dimensions of the original input image. This procedure resulted in a set of 1,460 attention tensors from each network, half veridical and half scrambled, with the same dimensions as the input images the networks were originally trained on.
Receiving networks were constructed by first pre-training them on the image discrimination task following the same training procedures described above. We then froze all weights in the network except those in the final multilayer perceptron (MLP) head (a module consisting of two linear layers, which computes the final classification decision). We further trained the network for 800 training epochs by providing the network with new training data in the form of 1,314 labeled attention tensors saved from the sending networks, half veridical and half scrambled. The receiving networks were then tested on an additional 146 attention inputs (half veridical, half scrambled). To avoid any possible bias from the original image classification task, receiving networks were pre-trained on one of the three image classifications only (e.g., classification A) and were then trained to judge attention data from networks trained on a different image classification task (e.g., B or C). At the end of training, the performance accuracy of the receiving network was saved. Since three receiving networks were trained, one for each of the three classification tasks, three accuracy scores were obtained. We then repeated the process 9 times, resulting in 27 accuracy scores. For all results displayed, each point represents a mean of 27 measurements.
@@
The experiment used a 2$\times$2 design in which the sender network could be a control network or an attention schema network and the receiver network could be a control network or an attention schema network. We hypothesized that networks with attention schemas should be better at modeling other agents’ attention than networks without schemas. Thus, a receiving network with an attention schema should be better at discriminating the attention data from sending networks. We also predicted that attention from networks with attention schemas should be easier to model than attention from networks without schemas. Thus, a receiving network, whether it does or does not have an attention schema, should be better at discriminating attention data from sending networks that have attention schemas.
@@
These effects were statistically significant. We performed a 2$\times$2 ANOVA and found that the main effect of the receiving network (receiving network does or does not have an attention schema) was statistically significant (F=10.17, p=0.0018). The main effect of the sending network (sending network does or does not have an attention schema) was also statistically significant (F=136.18, p=1.28$\times$10-20). The interaction between the two factors was also significant (F=4.99, p=0.027). The largest effect was whether the sending network contained an attention schema. Having an attention schema made the network more interpretable to other networks.
Figure 3: Figure 3. Performance of Schema and Control networks on image classification tasks. A. Mean accuracy on image classification task for schema versus control networks. Bars show 95% confidence interval about the mean. No significance difference was obtained. B. Accuracy of schema versus control networks on transfer learning (pretraining on one image classification task followed by transfer learning to a second image classification task). Bars show 95% confidence interval about the mean. No significance difference was obtained.
We also examined performance of networks on the visual categorization task. We did not have any specific hypothesis about whether a network with an attention schema would perform the visual categorization task better or worse than a network without an attention schema. Figure 3A shows the result. When analyzing all networks trained on the visual categorization task, comparing those with and without an attention schema, we saw no significant difference in performance (t = –0.34, p = 0.73). This result suggests that adding an attention schema to a network does not automatically make the network perform better due to a non-specific effect of adding computational complexity. This question of the specificity of the effect is revisited in Experiment 2. From the data thus far, it appears that the addition of the attention schema improved performance in a specific manner: it improved the ability of one network to judge the attention states of another network and it improved the ability of one network to produce attention states that are more easily judged by another.
@@
Experiment 1 showed that the presence of an attention schema in a network confers two specific advantages, at least in one kind of network and with a specific implementation of an attention schema. The first advantage is that when a network with an attention schema is pretrained on a task, that network is then better able to learn to make judgements about the attention states of other networks. Through initial training in which it predicts its own attention states, it gains an improved ability to model the attention states of others. The second advantage is that when a network contains an attention schema, its attention states become easier for other networks to interpret and categorize. In principle, both of these changes should be advantageous in joint tasks in which interpreting, categorizing, and predicting the behavior of others is useful. It may be of interest that the largest impact we observed here was that when a network contained an attention schema, its attention states were more easily recognized and categorized by other agents. It has been suggested that self-models such as attention schemas help to restructure networks, causing them to make their internal computations simpler and more regularized, and as a result could make those networks an easier target for others to predict and interpret [16]. The present results are consistent with that hypothesis.
@@
To check this possibility, we conducted transfer learning on the networks. First we pretrained them on the image classification task, replicating the procedure in Experiment 1. However, instead of then training the networks to classify attention (real versus shuffled), we trained the networks on another image classification task. Networks that were pretrained on one of the three image classification tasks (for example, task A) were retrained on a different one (for example, task B). We then asked whether transfer learning from one classification task to another was better for networks that contained an attention schema.
The results are shown in Figure 3B. Transfer learning was comparable between the two network architectures. No significant difference was found (t = –0.74, p = 0.46). These results show that schema networks and control networks perform similarly on basic discrimination tasks both before and after transfer learning.

### v2

# Testing Components of the Attention Schema Theory in Artificial Neural Networks

@@ Abstract
Growing evidence suggests that the brain uses an attention schema, or a simplified model of attention, to help control what it attends to. One proposed benefit of this model is to allow agents to model the attention states of other agents, and thus predict and interact with other agents. The effects of an attention schema may be examined in artificial agents. Although attention mechanisms in artificial agents are different from in biological brains, there may be some principles in common. In both cases, select features or representations are emphasized for better performance. Here, using neural networks with transformer attention mechanisms, we asked whether the addition of an attention schema affected the ability of agents to make judgements about and cooperate with each other. First, we found that an agent with an attention schema is better at categorizing the attention states of other agents (higher accuracy). Second, an agent with an attention schema develops a pattern of attention that is easier for other agents to categorize. Third, in a joint task where two agents must predict each other to paint a scene together, adding an attention schema improves performance. Finally, the performance improvements are not caused by a general increase in network complexity. Instead, improvement is specific to tasks involving judging, categorizing, or predicting the attention of other agents. These results support the hypothesis that an attention schema has computational properties beneficial to mutual interpretability and interactive behavior. We speculate that the same principles might pertain to biological attention and attention schemas in people.
@@ Introduction
Growing evidence suggests that the brain uses an attention schema, or a simplified model of attention, to monitor, predict, and help control attention [1-5]. The attention schema has also been hypothesized to play a prominent role in social cognition. It may be used to model the attention of others, affording better theory of mind, better predictions of the behavior of others, and therefore more effective social interaction [1,6-12]. One way to test hypotheses about a proposed computational mechanism in the brain, such as an attention schema, is to incorporate it into artificial agents, where its consequences can be tested in a controlled manner. In the present study, we tested whether adding an attention schema to an artificial agent improves its ability to make judgements about the attention of other agents, improves its ability to have its own attention judged by others, and improves its ability to coordinate with another agent in a joint task.
The biological mechanism of selective attention in the human brain, at least as it is understood, may be quite different from the kinds of attention that have been implemented in artificial agents. A great variety of attention mechanisms have been tested in machine learning, some more inspired by biological attention and some less. One particular method, associated with transformer architecture [13], has become extremely successful at helping artificial agents learn complex tasks, and now lies at the root of most large artificial intelligence (AI) agents such as modern chatbot technology. The extent of similarity or difference between transformer attention and biological, human attention has been controversial. Original accounts described the mechanism as similar, but later interpretations suggested that fundamental differences exist [14]. In the present study, we were inspired by proposed principles of the attention schema as derived from human cognitive neuroscience [1-12], and yet we tested these principles in an artificial transformer architecture. There are several reasons for this choice. First, we wished to examine the utility of the biologically-inspired attention schema when applied to the most common and successful mechanism of attention currently used in artificial systems. Though versions of attention schemas have been tested using other kinds of artificial attention [15-17], these previous studies have had limited impact and generated limited interest because they do not involve attention mechanisms that are mainstream in machine learning, and therefore have limited practical applicability.
Second, and most importantly, the specific mechanism by which attention is implemented should not change the underlying principles at work. Although the mechanisms may be quite different between biological and transformer attention, the deeper principles do share similarities, and it is these principles that matter most. Selective attention, whether in a human brain or an artificial transformer architecture, is fundamentally about selecting some features to have a greater influence over downstream information processing and therefore, ultimately, the behavior of the agent. The central concept in the attention schema theory is that, because of the importance of this selection, attention is one of the most important factors shaping behavior. Therefore, to better predict behavior, it is useful to build a predictive model of attention; and therefore, there is utility in building a predictive model of one’s own attention, and a social utility in building a model of the attentional states of others. This underlying concept is not tied to the specific implementation details of attention. For example, whether attention emerges from a biased competition mechanism of neuronal inhibition and excitation involving top-down and bottom-up signals, as for example some have theorized for the human brain, or whether attention emerges as a result of a transformer architecture involving multiple attention heads, as is often implemented in machine learning, the deeper concepts remain. A predictive model of an agent’s own attention mechanism, and of the attention mechanisms of other agents, should be of benefit, especially in social tasks that require fine-tuned prediction of the behavior of other agents. If these principles, inspired by the biological case, do apply to the very different implementation of transformer attention, then they may prove to be of great utility in building effective AI, and most especially, in shaping agents that have better social cooperative abilities.
Recently, versions of attention schemas have been incorporated into artificial neural networks to test whether the same principles of the attention schema inspired by data from humans might provide benefits in machine learning, and to determine whether lessons from the more controlled, artificial case might provide conceptual insights into human cognition [15-18]. Several studies suggest that an agent that controls movement of a spotlight of visual attention can do so more effectively when it includes a version of an attention schema to monitor and predict its attentional state [15-17]. These studies were the first to show the validity of the attention schema hypothesis in an artificial, computational context. However, they addressed only whether an attention schema aids an agent in controlling its own attention, while leaving untested the much more complex question of whether an attention schema aids an agent in making inferences about the attentional states of others or in the performance of interactive tasks with others.
@@
One recent study suggests that an attention schema may improve performance of artificial systems in cooperative tasks. Liu et al. [18] trained interacting groups of networks on visual tasks that required the agents to coordinate with each other. In that study, a transformer-style visual attention was used. In some conditions, networks were provided with an attention schema, in the form of a learned ability to predict the state of their own attention mechanism. With an attention schema, agents performed the social coordination tasks better. One interpretation is that when an agent is trained to predict its own attention states, it also becomes better at predicting the attention states of other agents in the same environment. In that hypothesis, learning to self-model leads to an improvement in the ability to model others and thus to an improvement in group task performance.
@@
A second recent study reported unexpected consequences for agents that build predictive self-models. An attention schema is a specific type of self-model, since it is a predictive model of the attentional state of an agent, but one can study the general impact on an agent of learning to predictively model any aspect of itself. Premakurma et al. [19] trained artificial networks on a variety of primary discrimination tasks, including both visual and linguistic tasks. In some conditions (the self-model conditions), in addition to the primary task, the agents were also trained on an auxiliary task in which they were rewarded for predicting the activation states of their own inner layers. The networks that were trained on the self-modeling task showed consistent changes. They learned to make themselves simpler, more regularized, and more parameter-efficient. These gains were found across all types of networks studied. It was hypothesized that when a network is given the auxiliary task of predicting its own internal activity states, it learns to make those internal states simpler and more regular. The simpler the internal states, the easier they are to predict. Therefore, as the network learns to predict itself, it learns to simplify itself. According to that hypothesis, learning a self-model restructures the self, regularizing internal processes, and making internal states an easier target of predictive modeling.
@@
Based on these prior findings, we proposed the following two hypotheses about the relationship between self-models and social behavior in artificial agents. First, when an agent learns to self-model, such as when it learns an attention schema that models its own attention state, it can transfer some of that learning toward predictively modeling other agents. Predictive modeling of other agents then becomes a useful part of interacting with them. Second, when an agent learns to self-model, it restructures itself to be more regularized and more predictable, and therefore the agent becomes a more effective target for being predictively modeled by other agents. Both changes – becoming better at modeling others, and becoming a more transparent target for being modeled by others – should render the agent a more effective member of a cooperative group.
@@
The purpose of the present study was to test these hypotheses about how attention schemas might influence multi-agent interaction in artificial agents using a transformer style of attention. We conducted three related experiments. Experiment 1 asked: Does an agent that has an attention schema that models its own attention also have an improved ability to model the attention states of other agents? Experiment 2 tested an alternative explanation: Are the performance benefits of an attention schema merely a non-specific effect of added network complexity? Finally, Experiment 3 addressed the question of multi-agent cooperation: Do agents with attention schemas perform better on a simple joint task?
@@
The purpose of experiment 1 was to train artificial agents with and without attention schemas. The agents used a transformer type of attention. The attention schema, or learned predictive model of its own attention state, was adapted to the transformer type of attention mechanism, as described below. We tested two hypotheses. Hypothesis 1 was that when an agent has an attention schema, it will be better at learning to make categorical judgements about the attention of other agents. Hypothesis 2 was that when an agent has an attention schema, other agents will be better at learning to make categorical judgements about its attention.
@@
Agents were first trained on a visual categorization task (e.g. categorize a picture as showing a golf ball or a garbage truck). We then froze the weights in all layers except a final block, and trained the agents on a new task, the “attention judgement” task. The task allowed us to test the ability of one agent to make categorical determinations about the attention states of another agent. The attention judgment task was modified from Ziman et al. [20], who demonstrated the ability of people to assess the attention states of others. In the human version of the task, participants looked at a spotlight of attention moving around a picture and judged whether the spotlight represented real, overt attention (based on eye movements recorded previously from a real person), or whether it represented a randomly scrambled version of attention (the recorded attention locations presented in a random order). Since people were able to perform above chance, the human brain must contain information about the typical or predicted patterns of attention and must be able to recognize when those patterns have been violated. Thus, humans demonstrate an implicit predictive model of attention. It is not simply the case that they can learn to distinguish one arbitrary type of pattern from another – rather, they know, without any feedback, which patterns are typical of real human attention and which ones are not.
@@
To modify the task for the present study such that it could be applied to a transformer type of attention, we assigned agents to one of two roles: either the receiver or the sender. If the agent was assigned the sender role, it performed its originally trained visual categorization task. The attention activation values, for each trial of the task, were saved in the form of a multidimensional tensor. In one version, the tensor was left in its correct, original form. In another version, the attention values were randomly scrambled along one dimension of the tensor.
@@
If the agent was assigned the receiver role, then it was first trained on the image categorization task; then its weights were frozen except in the final block of layers; and then it was trained on the attention judgement task. On each trial of the attention judgement task, instead of receiving a picture as an input, the agent was provided with an attention tensor; and instead of being rewarded for categorizing the picture, it was rewarded for categorizing the tensor as either a valid example of attention (recorded from the performance of a sender agent) or as an invalid example (constructed by randomly modifying a valid tensor). The task therefore tested the ability of the receiver agent to learn to recognize violations of predictable patterns in the attention behavior of the sender.
@@
The experiment used a $\times2$ design: the sender agent could be trained with or without an attention schema (a learned ability to predict its own attention states), and the receiver agent could be trained with or without an attention schema. We proposed the following hypotheses.
@@
Hypothesis 2: If the sender agent was trained with an attention schema, then the receiver agent will be able to learn the attention judgement task better. Such a result would suggest that when an agent contains an attention schema, it is, itself, an easier target for other agents to build predictive models of it. It would suggest that having an attention schema adjusts or restructures the agent’s own attention, to make it more easily interpretable to others.
@@
We trained networks on a visual categorization task. Given an image as input (for example, a picture of a golf ball on a lawn), the network was trained to arrive at a dichotomous categorization (for example, golf ball versus garbage truck). All networks and task environments were created using PyTorch [21] and all experiment code and data are publicly available at github.com/kathrynfarrell/attention_schemas_in_anns. We used an image-adapted implementation of the Attention Schema Neural Network architecture proposed by Liu et al. [18], which utilized a transformer-based attention mechanism. Figure 1 shows the architecture, with 1A showing the control architecture (control network: attention but no attention schema) and 1B showing the architecture that includes an added attention schema (attention schema network). It is important to note that the task of categorizing one type of picture from another is not necessarily attentionally demanding for humans, but relies on the transformer attention mechanism in the artificial agent. To solve the task, the agent must be able to select features in the picture to have greater impact, and it is the transformer encoder that performs that function. Thus, this is an appropriate task for testing the attention hypotheses as applied to these artificial agents.
@@
Figure 1: Figure 1: The Attention-Schematic Neural Network (ASNN) architecture. A. The control model without an attention schema. B. The model with an attention schema, modified from Liu et al. [18]. The model is based around a standard vision transformer (ViT) block (“Transformer Encoder”), which generates attention scores and applies them to the network’s representation of the image (Values), to produce a representation of the image with the attention applied (“applied attention”). Conceptually, the attention scores can be thought of as a heatmap of where to focus and the Applied Attention can be thought of as the heatmap overlaid onto the image. The applied attention is then passed into a Recurrent Neural Network (RNN) module with learnable parameters. The output of this module ($R_{out}$”) is a precursor to both the network’s final attention prediction and its dynamic selection of attention scores, both of which are produced by several linear layers with learnable parameters (the Predictive Layers and the Decision Control Layers, respectively). For the attention-revision computations, attention suppression or activation is achieved through processes beginning in the Decision Control Layers. Linear “suppressor” and “activator” layers in the Decision Control produce logit pairs, which are binarized using a Gumbel-Softmax distribution, then combined via element-wise multiplication. The binary mask resulting from that multiplication is then applied to the original attention scores (by element-wise multiplication) yielding the final attention—a sparser version of the original attention scores where some of the scores have been selectively “shut off” or reduced to zero. The final attention is then applied to the original image representation (Values), and the result is passed to the final MLP head, which classifies the image as belonging to a particular category. This MLP head is the only part of the network that continues updating during transfer learning (specifically, the final multi-layer perceptron in the MLP head), while the rest of the network remains frozen.
@@
For details on how the attention schema mechanism works, see Liu et al. [18]. Briefly, the objectives of the network with the schema are twofold: learning to model and refine its own attention and learning to complete a task (e.g. classifying images). The schema network uses its applied attention in service of both goals: modeling its own attention (by predicting its final attention), and further editing the attention to its final refined state. To accomplish this, the schema network uses a transformer block with six attention heads, and passes the applied attention into a recurrent neural network module with learnable parameters. The use of multiple attention heads is standard in transformer architectures. The duplication allows for better performance. Again, although this architecture is very different from the implementation of attention in the human brain, it is the most effective way to test whether the general principles of an attention schema apply to standard implementations in machine learning. The output of this module ($R_{out}$) is further processed and used to model attention; specifically, it is passed into the Predictive Layers, which output the predicted attention. $R_{out}$ is also further processed and used to refine attention; it is passed into the Decision Control Layers which produce the final refined attention (with the Decision Control Layers employing activator units, suppressor units, and categorical reparameterization with gumbel-softmax). In this way, manipulations of the initial applied attention are used to generate both the prediction of the final attention as well as the final attention itself. Schema networks are trained on their combined ability to accurately complete the main task (e.g. classifying an image) and predict their own attention.
@@
To produce an attention tensor, we stacked the scores from the three attention heads into a 257 token $\times$ 257 token $\times$ 3 attention head tensor (257 tokens in the first two dimensions resulting from the 16 $\times$ 16 image patches from the original input image plus one class token). From these tensors, we created the veridical and scrambled attention data that would be sent to the receiving networks for discrimination. The veridical attention tensors were created by reducing the size of the tensors from 257 $\times$ 257 $\times$ 3 to 256 $\times$ 256 $\times$ 3, matching the size of the original input image. For the scrambled attention, we randomly permuted each tensor along its last dimension, which randomly shuffled the attention scores within the data from each attention head. Then, we reduced the resulting tensor to 256 $\times$ 256 $\times$ 3, again matching the dimensions of the original input image. This procedure resulted in a set of 1460 attention tensors from each network, half veridical and half scrambled, with the same dimensions as the input images the networks were originally trained on.
Receiving networks were constructed by first pre-training them on the image discrimination task following the same procedures described above. We then froze all weights in the network except those in the final multilayer perceptron (MLP) head (a module consisting of two linear layers, which computes the final classification decision). We further trained the network for 800 training epochs by providing the network with new training data in the form of 1,314 labeled attention tensors saved from the sending networks, half veridical and half scrambled. The receiving networks were then tested on an additional 146 attention inputs (half veridical, half scrambled). To avoid any possible bias from the original image classification task, receiving networks were pre-trained on one of the three image classifications only (e.g., classification A) and were then trained to judge attention data from networks trained on a different image classification task (e.g., B or C). At the end of training, the performance accuracy of the receiving network was saved. Since three receiving networks were trained, one for each of the three classification tasks, three accuracy scores were obtained. We then repeated the process 9 times, resulting in 27 accuracy scores. For all results displayed, each point represents a mean of 27 measurements.
@@
Figure 2 shows the results for experiment 1. Performance on the attention discrimination task was well above chance (mean = 80.87% correct, SEM = 2.38) when both the sending network and the receiving network contained an attention schema. Performance dropped when the sending network contained an attention schema but the receiving network did not (mean = 70.61, SEM = 2.37). Performance dropped further when the sending network did not have an attention schema, but the receiving network did (mean = 54.58, SEM = 1.25). Finally, performance was lowest when neither sending nor receiving network contained an attention schema (mean = 52.78, SEM = 1.21). These results support both hypotheses. With respect to hypothesis 1: When the receiving network contained an attention schema, it had an advantage in learning to categorize the attention data from the sending network. With respect to hypothesis 2: When the sending network contained an attention schema, the network was in some manner altered such that its attention data was easier for the receiving network to categorize.
@@
These effects were statistically significant. We performed a 2 $\times$ 2 ANOVA and found that the main effect of the receiving network (receiving network does or does not have an attention schema) was statistically significant (F=10.17, p=0.0018, $\eta^{2}$p=0.09, with a post-hoc power analysis indicating statistical power of 0.77). The main effect of the sending network (sending network does or does not have an attention schema) was also statistically significant (F=136.18, p=1.28 * 10-20, $\eta^{2}$p=0.57). The interaction between the two factors was also significant (F=4.99, p=0.027, $\eta^{2}$p=0.05). The largest effect was whether the sending network contained an attention schema. Having an attention schema made the network more interpretable to other networks.
Figure 3: Figure 3. Performance of Schema and Control networks on image classification tasks. A. Mean accuracy on image classification task for schema versus control networks, in experiment 1. Bars show 95% confidence interval. No significance difference was found. B. Results for experiment 2. Mean accuracy of schema versus control networks on transfer learning (pretraining on one image classification task followed by transfer learning to a second image classification task). Bars show 95% confidence interval. No significance difference was found.
We also examined performance of the networks on the visual categorization task. We did not have any specific hypothesis about whether a network with an attention schema would perform the visual categorization task better or worse than a network without an attention schema. One possibility is that the attention schema would improve the agent’s ability to refine its own attention, and thus improve its performance on the categorization task. Another possibility is that the presence of the attention schema would provide benefits only for the more “social” aspects of performance, categorizing the attentional states of other agents. Figure 3A shows the result. When analyzing all networks trained on the visual categorization task, comparing those with and without an attention schema, we saw no significant difference in performance (t=-0.34, p=0.73, d=0.09, post-hoc power=0.06). While acknowledging that null results must be treated with caution, we interpret this result as suggesting that adding an attention schema to a network does not make the network automatically perform better whether due to more refined attention or due to a non-specific effect of adding computational complexity. This question of the specificity of the effect is revisited in Experiment 2. From the data thus far, it appears that the addition of the attention schema improved performance in a specific manner: it improved the ability of one network to judge the attention states of another network, and it improved the ability of one network to produce attention states that are more easily judged by another. In this sense, it improved more “social” aspects of performance.
@@
Experiment 1 showed that the presence of an attention schema in a network (the trained ability to predict its own attention states) confers two specific advantages, at least in one kind of network with a specific implementation of attention and of an attention schema. The first advantage is that when a network with an attention schema is pretrained on a task, that network is then better able to learn to make judgements about the attention states of other networks. Through initial training in which it predicts its own attention states, it gains an improved ability to model the attention states of others. The second advantage is that when a network contains an attention schema, its attention states become easier for other networks to interpret and categorize. In principle, both of these changes should be advantageous in joint tasks in which interpreting, categorizing, and predicting the behavior of others is useful. It may be of interest that the largest impact we observed here was that when a network contained an attention schema, its attention states were more easily recognized and categorized by other agents. It has been suggested that an important feature of social interaction, whether in people or in simple artificial agents, is that self-models (such as attention schemas) help to restructure networks, causing them to make their internal computations simpler and more regularized, thus making those networks an easier target for others to predict and interpret [19]. The present results are consistent with that hypothesis.
@@
To check this possibility, we examined the networks’ performance in transfer learning. First, we pretrained them on the image classification task, replicating the procedure in experiment 1. However, instead of then training the networks to classify attention (real versus shuffled), we trained the networks on another image classification task. Networks that were pretrained on one of the three image classification tasks (for example, task A) were retrained on a different one (for example, task B). We then asked whether transfer learning from one classification task to another was better for networks that contained an attention schema.
The results are shown in Figure 3B. Transfer learning was comparable between the two network architectures. No significant difference was found (t=-0.74, p=0.46, d=0.20, post-hoc power=0.111). While null results should be considered with caution, these results generally suggest to us that schema networks and control networks perform similarly on basic discrimination tasks both before and after transfer learning.